use the method of substitution to solve the system of linear equation
a. 2x-y=2
b. 7x-3y=10

Answer by mananth(16949)

2 x + - y = 2 .............1
- y = 2 + -2 x
/ -1 =
y = -2 + 2 x
7 x + -3 y = 10 .....................2
Substitute y in (2)
7 x -3 ( -2 + 2 x ) = 10
7 x + 6 + -6 x = 10
7 x -6 x = 4
1 x = 4
/ 1
x= 4
Plug the value of x in (1)
2 x - y = 2
2 * 4 - y = 2
8 + - y = 2
-1 y = -6
/ -1
y= 6
m.ananth@hotmail.ca